# Bridging Earth and Space: A Survey on HAPS for Non-Terrestrial Networks

### v1

@@ V-A Channel Modeling
A multi-state statistical channel model for HAPS communication in the Ka-band was proposed in [122] to account for tropospheric weather effects and large-area user mobility. The model uses three fading states—LoS, moderate shadowing, and deep shadowing—represented by a Markov chain. It also provides the probability density function of the received signal power, a method for generating time-series, and evaluates BER performance using the proposed model.
The authors in [123] showed Ka-band rain attenuation is about 10× higher than in C-band and strongly location-dependent, and presented an analytical model based on the International Telecommunication Union Radiocommunication Sector (ITU-R) empirical model to predict Earth–space attenuation for fully and partially rain-affected paths. For a 1-km cell, attenuation reaches 43 dB at 0.01% availability and roughly doubles as cell size increases. Site diversity was evaluated as a rain-fade mitigation technique for HAPS-Earth links using the ITU-R model with three Earth stations at different rainfall rates to predict rain attenuation statistics [124]. Results show that site diversity raises link availability from 97% to 99.9%, as indicated by improvements in carrier-to-noise ratio.
@@
The study in [125] shows that HAPS mobile channels invalidate the terrestrial plane-wave assumption, requiring consideration of both elevation and azimuth spreads. Large delay differences, Doppler shifts, and strong spatial correlation further complicate analysis. By characterizing dispersion across time, frequency, and space, the authors develop a wideband tapped-delay-line HAPS channel model with environment-specific parameters to support simulation and verification of broadband HAPS systems.
The study in [121] examined coherence bandwidth and RMS delay spread of satellite channels at Ku- and X-bands across four propagation scenarios. It found lower delay spread in woodland areas, abrupt location-dependent variations in rural environments, and urban delay spread marginally lower than suburban due to fewer, stronger multipath components with smaller relative delays.
FSO communication provides several benefits, e.g., in terms of bandwidth, data rate, and licensing costs, compared to traditional technologies, as described in Sec. IV-B. However, it faces significant challenges due to its dependency on precise transmitter-receiver alignment and the physical properties of the medium. Along these lines, the authors in [126] focused on optimizing internal parameters for FSO systems such as wavelength, detector type, beam divergence, modulation, and power to maximize link range and availability.
V-B Resource Management
@@
Given the recent relevance of HAPS in wireless communication, resource allocation (RA) has become one of the key areas of investigation, including the positioning of the aerial tier units within the network. While multiple parameters must be optimized to address the RA problem comprehensively, more practical solutions operate under specific circumstances and constraints, defining optimization objectives accordingly. In this subsection, we will consider various system models, from single-tier (based on HAPS) to multi-tier (also known as space-air-ground integrated networks) architectures.
V-B1 Single-tier architecture
@@
In [127], the authors maximize the number of UEs receiving multicast in an orthogonal frequency division multiple access (OFDMA) frame under QoS constraints by decomposing the problem into two linear programs solved iteratively. Sub-problem 1 selects UEs and assigns frequency–time slots via three Lagrangian-relaxation methods. Sub-problem 2 allocates power, modeled as a continuous knapsack, solved optimally with a greedy algorithm. The scenario assumes a single HAPS serving ground user equipments in a defined area.
@@
In a multi-platform setup with several HAPS serving an area, [128] exploits fixed user dishes with narrow beamwidths steered to their main HAPS to boost bandwidth efficiency. Two cases are studied—single-beam per HAPS and multi-beam per platform—and results show higher capacity when multiple HAPS share the same spectrum, with gains that scale as the number of users grows.
In multi-cell settings, both inter- and intra-beam interference are significant. In integrated networks, overlapping coverage between terrestrial base stations and HAPS-based BSs—especially with co- or adjacent-channel operation—induces inter-system interference, making effective interference coordination essential for future integration (see Sec. V-H).
@@
In [129], a dynamic spatial–time–frequency coordination method between HAPS-mounted BSs and TBSs is proposed. Two service modes are considered: direct HAPSGUE links and indirect access where HAPS serve TBSs that, in turn, serve GUEs. UEs associate to HAPS (direct or indirect) or TBSs based on maximum received power. The scheme distributes traffic load and employs an ergodic search over allocation patterns across HAPS and terrestrial networks. Simulations show the best average per-cell throughput for user groups at any distance from HAPS.
Unlike the single-HAPS model in [129], [26] studies a HAPS network via stochastic geometry, modeling TBSs and HAPS as Poisson point processes and deriving a closed-form coverage probability for spectrum sharing. Results show HAPS cause minimal interference to TBSs, supporting spectrum sharing in integrated networks.
V-B2 Multi-tier architecture
@@
The study in [131] focused on RA optimization within a multi-tier architecture where LEO satellites provide wide-area coverage, while HAPS and full-duplex terrestrial relays cover hot-spot regions to enhance capacity. Additionally, GEO satellites were considered as routing nodes. The authors designed a hierarchical resource management framework for this system, and proposed an optimization method to maximize LEO-to-ground and HAPS-to-ground downlink throughput based on predictions derived from the regular orbit information of LEO satellites. In the proposed architecture, mobile UEs in hot-spot regions connect via HAPS or terrestrial relays, while other areas use direct LEO links. This approach boosts system capacity in hot-spots and lowers handover rates. Different frequency bands are allocated for space/air-to-ground, terrestrial, space-to-air, and inter-satellite links. Capacity analysis and simulations show that the multi-tier design greatly outperforms satellite-only systems. Additionally, dynamic radio resource optimization improves the average throughput of HAPS and LEO beams by over 10%. In this scenario, spectrum sharing is required if HAPS and satellites are designed to transmit on the same or adjacent frequency bands, or when deploying HAPS in a region already covered by a separate satellite system. A LEO-HAPS spectrum sharing technique under non-ideal spectrum sensing was studied in [132], where a multi-beam LEO served as the primary system and a single-beam HAPS supported GUEs without direct LEO access. A dynamic spectrum and power allocation strategy was proposed to maximize HAPS-GUE throughput with minimal QoS degradation for LEO-associated GUEs. Results showed that HAPS interference to LEO-associated GUEs is negligible and affects only a small fraction.
@@
RA for a multi-tier cognitive satellite-HAPS network with non-orthogonal multiple access (NOMA) was studied in [133]. The model included a primary satellite and a secondary HAPS network sharing spectrum via underlay cognitive radio to serve terrestrial users. The sum-rate maximization problem for the secondary network was addressed under NOMA group connectivity limits, power constraints, and QoS requirements. The proposed joint RA algorithm (sub-channel and power allocation) outperformed orthogonal multiple access, especially as user or sub-channel numbers increased.
@@
Introducing smaller aerial components, such as LAPSs like drones, into the network architecture adds an additional degree of freedom in deployment planning. While this approach can further improve coverage and system capacity, it also complicates RA due to the need for dynamic load balancing and 3D placement of mobile network nodes. An aerial heterogeneous cellular network with HAPS-based macrocells and LAPS-based femtocells was studied in [134]. Using stochastic geometry, the authors analyzed cell loads, outage probability, optimal LAPS placement, and downlink signal to interference plus noise ratio (SINR). Results showed that users associated with an LAPS BS experienced much lower outage probability, and the proposed LAPS positioning algorithm further reduced outages.
@@
The study in [135] investigated joint RA and UAV 3D-location optimization in a cluster-NOMA network with a HAPS macro BS and multiple UAV small BSs. The goal was to maximize IoT uplink sum rate by decomposing the problem into spectrum/power allocation and UAV positioning (horizontal and altitude), solved iteratively. Results showed improved uplink sum rate and spectrum efficiency over benchmark RA schemes.
As previously mentioned, the SAGIN architecture is likely the most promising option for future communication networks, including 6G. Therefore, it is crucial to clarify the features that enable the coexistence and collaboration of HAPS and terrestrial systems within this integrated framework. In [136], a RA framework for an integrated satellite-HAPS-TBS network was proposed to improve GUE downlink throughput under power, access, and backhauling constraints. Two problems were formulated: a short-term problem optimizing access, backhauling, and power with a fixed HAPS location, and a long-term problem optimizing HAPS placement based on average user distribution.
@@
The studied system model comprised three tiers: a satellite in the space tier, multiple HAPS in the air tier, and several TBSs in the ground tier, with gateway feeder stations for backhauling. Based on OFDMA, spectrum was divided into resource blocks, with each user served by one BS using one RB, ensuring no intra-cell interference but allowing inter-cell interference among TBSs. No inter-cell interference was assumed for satellites and HAPS due to distinct frequency sets. Simulations showed a notable average data rate gain compared to two baselines: (i) joint optimization with uniform power, and (ii) optimized HAPS placement with random access/backhaul under uniform power.
@@
From the literature, it is clear that aerial network mobility boosts capacity and flexibility but complicates optimization due to 3D positioning, dynamic load balancing, and spectrum diversity. Effective interference management is vital for stability, while a well-designed RA strategy can mitigate inter-tier interference and enable smooth integration with terrestrial networks.
V-C User Assignment and Cell Association
@@
Unlike stationary TBSs, HAPS offer flexible deployment that can be optimized to user distribution. The study in [137] examined HAPS deployment for maximizing network transmission rate while meeting service requirements, where each GUE connects to the HAPS offering the best rate. A game-theoretic, self-organized optimization using a restricted spatial adaptive play algorithm was proposed and shown to converge in simulations. Although simplified compared to full SAGIN architectures, the results highlighted the potential of HAPS for self-organized deployment optimization.
@@
To meet 6G data rate demands, [138] studied user selection in a downlink NOMA-based multi-user HAPS system. A joint user selection and correlation-based pairing algorithm was proposed to maximize beamforming gain, along with a polar-cap codebook design for limited-feedback HAPS channels under Rician fading. Simulations showed superior data rates over existing NOMA schemes. However, the model did not consider integration with terrestrial/satellite networks or support for aerial users, which are essential for 6G.
@@
The user association problem in the context of computational task offloading was studied in [139]. The authors considered a disaster scenario where TBSs have been destroyed, and two-tier UEs, including both GUEs and UAVs, with limited computational capabilities offload disaster information to HAPS and/or LEO satellite servers for further processing. The study aimed to jointly optimize uplink fronthaul capacity and computation performance by modeling the problem as a weighted 3-uniform hypergraph, where each hyperedge weight represents the combined fronthaul uplink and computing rates. A local search-based algorithm was developed to find a near-optimal 3D hypergraph matching with maximum total weight.
The model considered a LEO satellite with an enhanced MEC server providing backhaul to multiple HAPS, which cover disaster areas and offload computation, while UEs use orthogonal sub-channels within shared Ka-band RBs, avoiding uplink interference. In simulations with 6 HAPS and 20 UAVs over a 0.25 km² disaster area, the proposed hypergraph search algorithm achieved significant system sum rate gains over greedy search, showing strong practical potential.
In [140], HAPS-mounted BSs were studied as backup transmitters in areas with strong terrestrial coverage. The user association problem between a TBS and a HAPS was addressed under outdated channel state information (CSI), with a DRL-based algorithm designed to maximize downlink sum-rate without requiring HAPS-user CSI. The TBS-user CSI was updated only for users served in the previous slot, and scenarios with imperfect CSI were also considered. Simulations confirmed the algorithm’s near-optimality, stability, and scalability despite CSI limitations.
@@
A vertical heterogeneous network (VHetNet) comprising one GEO satellite, one HAPS, and several TBSs was examined in [141]. In this setup, both the HAPS and TBSs are equipped with multiple antennas, enabling them to simultaneously serve multiple GUEs through RF links. The connection between the satellite and the HAPS is established via an FSO link. In this scenario, when a GUE is served by the HAPS, the required data is first transmitted from the satellite to the HAPS, and then relayed to the GUE. The system model employs a space division multiplexing scheme, where all RF links share the same central frequency. Consequently, the performance of the proposed architecture is highly dependent on the intra-cell and inter-cell interference among TBSs, intra-cell interference within the HAPS, and the interference between HAPS and TBSs. The study formulated a joint optimization problem for associating GUEs with either TBSs or the HAPS and designing corresponding beamforming vectors, aiming to maximize sum-rate under constraints of HAPS payload connectivity, FSO backhaul capacity, and power limits of HAPS/TBSs. An iterative algorithm optimized parameters step by step, and simulations demonstrated that integrated satellite-HAPS-terrestrial networks can significantly enhance the sum-rate.
@@
As heterogeneous integrated networks face load imbalance from uneven user distribution and mobility, [142] proposed a priority-based load-balancing scheme for a two-layer downlink SAGIN using network slicing to mitigate underutilization and overload across cells. The proposed two-layer radio access network (RAN) includes TBSs in the terrestrial layer (priority for user association) and UAVs/LEO satellites in the non-terrestrial layer. Overloaded TBSs offload traffic to the non-terrestrial layer via RAN slicing into high-throughput, low-delay, and wide-coverage slices, with lower-priority slices preferentially offloaded. The optimization problem jointly considered throughput, delay, coverage, user association, resource allocation, and UAV deployment. Results showed improved slice throughput and average delay. While HAPS were not included, the study tackled the critical load-balancing challenge in SAGIN.
In [143], load balancing for 5G integrated satellite-terrestrial networks was studied using multiple radio access technologys. The approach handled both intra- and inter-RAT UE offloading from overloaded cells by introducing the radio resource utilization ratio as a common load metric, enabling efficient traffic redistribution without overloading target cells. Simulations showed improved throughput and QoS. While limited to a GEO satellite, the metric can be generalized to other NTN components, including HAPS.
In [144], load balancing was studied as a joint radio resource management problem—covering bandwidth allocation, UE association, and power control—in an integrated TN-NTN rural downlink network with multiple macro TBSs and LEO satellites. The proposed framework maximized the sum of log throughput across all UEs, improving overall network capacity.
The framework tackled the optimization sequentially—first user association and bandwidth allocation under fixed power, then transmit power with the first two fixed. Simulations showed notable throughput gains over the maximum reference signal received power (RSRP) baseline and confirmed that integrating TNs and NTNs in rural areas can greatly reduce out-of-coverage UEs.
In [109], a framework for bandwidth allocation, user association, and power control was proposed to balance capacity and energy consumption by adapting to traffic conditions. The approach improved mean throughput during high-traffic periods and reduced power consumption during low-traffic periods.
@@
Based on the reviewed studies, user association in SAGINs with HAPS-mounted BSs integrated with TNs still requires further research. Future work should address scenarios with a realistically large number of mobile ground and aerial UEs, while accounting for practical constraints from interference management and load balancing.
V-D Multi-Link Connectivity

### v2

@@ VI-A Channel Modeling
A multi-state statistical channel model for HAPS communication in the Ka-band was proposed by Zhao et al. [248] to account for tropospheric weather effects and large-area user mobility. The model uses three fading states—LoS, moderate shadowing, and deep shadowing—represented by a Markov chain. It also provides the probability density function of the received signal power, a method for generating time-series, and evaluates BER performance using the proposed model.
@@
Guan et al. [83] show that HAPS mobile channels invalidate the terrestrial plane-wave assumption, requiring consideration of both elevation and azimuth spreads. Large delay differences, Doppler shifts, and strong spatial correlation further complicate analysis. By characterizing dispersion across time, frequency, and space, the authors develop a wideband tapped-delay-line HAPS channel model with environment-specific parameters to support simulation and verification of broadband HAPS systems.
VI-B Resource Management
@@
Given the recent relevance of HAPS in wireless communication, RA has become one of the key areas of investigation, including the positioning of the aerial tier units within the network. While multiple parameters must be optimized to address the RA problem comprehensively, more practical solutions operate under specific circumstances and constraints, defining optimization objectives accordingly. In this subsection, we will consider various system models, from single-tier (based on HAPS) to multi-tier (also known as space-air-ground integrated networks) architectures.
VI-B1 Single-tier architecture
@@
Ibrahim and Alfa [100] maximize the number of UEs receiving multicast session transmissions in an orthogonal frequency division multiple access (OFDMA) frame under QoS constraints by decomposing the problem into two linear programs solved iteratively. Sub-problem 1 selects UEs and assigns frequency–time slots via three Lagrangian-relaxation methods. Sub-problem 2 allocates power, modeled as a continuous knapsack, solved optimally with a greedy algorithm. The scenario assumes a single HAPS serving GUEs in a defined area.
@@
In a multi-platform setup with several HAPS serving an area, Grace et al. [80] exploits fixed user dishes with narrow beamwidths steered to their main HAPS to boost bandwidth efficiency. Two cases are studied—single-beam per HAPS and multi-beam per platform—and results show higher capacity when multiple HAPS share the same spectrum, with gains that scale as the number of users grows.
@@
In multi-cell settings, both inter- and intra-beam interference are significant. In integrated networks, overlapping coverage between terrestrial base stations and HAPS-based BSs—especially with co- or adjacent-channel operation—induces inter-system interference, making effective interference coordination essential for future integration (see Sec. VI-H) In this context, Liu et al. [140] proposed a dynamic spatial–time–frequency coordination method between HAPS-mounted BSs and TBSs. Two service modes are considered: direct HAPSGUE links and indirect access where HAPS serve TBSs that, in turn, serve GUEs. UEs associate to HAPS (direct or indirect) or TBSs based on maximum received power. The scheme distributes traffic load and employs an ergodic search over allocation patterns across HAPS and terrestrial networks. Simulations show the best average per-cell throughput for user groups at any distance from HAPS. Unlike the single-HAPS model in [140], Wei et al. [227] studies a HAPS network via stochastic geometry, modeling TBSs and HAPS as Poisson point processes and deriving a closed-form coverage probability for spectrum sharing. Results show HAPS cause minimal interference to TBSs, supporting spectrum sharing in integrated networks.
VI-B2 Multi-tier architecture
@@
Li et al. [134] focused on RA optimization within a multi-tier architecture where LEO satellites provide wide-area coverage, while HAPS and full-duplex terrestrial relays cover hot-spot regions to enhance capacity. Additionally, GEO satellites were considered as routing nodes. The authors designed a hierarchical resource management framework for this system, and proposed an optimization method to maximize LEO-to-ground and HAPS-to-ground downlink throughput based on predictions derived from the regular orbit information of LEO satellites. In the proposed architecture, mobile UEs in hot-spot regions connect via HAPS or terrestrial relays, while other areas use direct LEO links. This approach boosts system capacity in hot-spots and lowers handover rates. Different frequency bands are allocated for space/air-to-ground, terrestrial, space-to-air, and inter-satellite links. Capacity analysis and simulations show that the multi-tier design greatly outperforms satellite-only systems. Additionally, dynamic radio resource optimization improves the average throughput of HAPS and LEO beams by over 10%. In this scenario, spectrum sharing is required if HAPS and satellites are designed to transmit on the same or adjacent frequency bands, or when deploying HAPS in a region already covered by a separate satellite system.
Meanwhile, a LEO-HAPS spectrum sharing technique under non-ideal spectrum sensing was studied by Wang et al. in [222], where a multi-beam LEO served as the primary system and a single-beam HAPS supported GUEs without direct LEO access. A dynamic spectrum and power allocation strategy was proposed to maximize HAPS-GUE throughput with minimal QoS degradation for LEO-associated GUEs. Results showed that HAPS interference to LEO-associated GUEs is negligible and affects only a small fraction.
@@
Later, RA for a multi-tier cognitive satellite-HAPS network with non-orthogonal multiple access (NOMA) was studied by Liu et al. [138]. The model included a primary satellite and a secondary HAPS network sharing spectrum via underlay cognitive radio to serve terrestrial users. The sum-rate maximization problem for the secondary network was addressed under NOMA group connectivity limits, power constraints, and QoS requirements. The proposed joint RA algorithm (sub-channel and power allocation) outperformed orthogonal multiple access, especially as user or sub-channel numbers increased.
@@
Introducing smaller aerial components, such as LAPSs like drones, into the network architecture adds an additional degree of freedom in deployment planning. While this approach can further improve coverage and system capacity, it also complicates RA due to the need for dynamic load balancing and 3D placement of mobile network nodes. An aerial heterogeneous cellular network with HAPS-based macrocells and LAPS-based femtocells was studied by Helmy and Arslan [93]. Using stochastic geometry, the authors analyzed cell loads, outage probability, optimal LAPS placement, and downlink signal to interference plus noise ratio (SINR). Results showed that users associated with an LAPS BS experienced much lower outage probability, and the proposed LAPS positioning algorithm further reduced outages.
@@
Later, Qin et al. [175] investigated joint RA and UAV 3D-location optimization in a cluster-NOMA network with a HAPS macro BS and multiple UAV small BSs. The goal was to maximize IoT uplink sum rate by decomposing the problem into spectrum/power allocation and UAV positioning (horizontal and altitude), solved iteratively. Results showed improved uplink sum rate and spectrum efficiency over benchmark RA schemes.
@@
As previously mentioned, the SAGIN architecture is likely the most promising option for future communication networks, including 6G. Therefore, it is crucial to clarify the features that enable the coexistence and collaboration of HAPS and terrestrial systems within this integrated framework. In [24], a RA framework for an integrated satellite-HAPS-TBS network was proposed to improve GUE downlink throughput under power, access, and backhauling constraints. Two problems were formulated: a short-term problem optimizing access, backhauling, and power with a fixed HAPS location, and a long-term problem optimizing HAPS placement based on average user distribution. The studied system model comprised three tiers: a satellite in the space tier, multiple HAPS in the air tier, and several TBSs in the ground tier, with gateway feeder stations for backhauling. Based on OFDMA, spectrum was divided into resource blocks, with each user served by one BS using one RB, ensuring no intra-cell interference but allowing inter-cell interference among TBSs. No inter-cell interference was assumed for satellites and HAPS due to distinct frequency sets. Simulations showed a notable average data rate gain compared to two baselines: (i) joint optimization with uniform power, and (ii) optimized HAPS placement with random access/backhaul under uniform power.
@@
From the literature, it is clear that aerial network mobility boosts capacity and flexibility but complicates optimization due to 3D positioning, dynamic load balancing, and spectrum diversity. Effective interference management is vital for stability, while a well-designed RA strategy can mitigate inter-tier interference and enable smooth integration with TN.
VI-C User Assignment and Cell Association
@@
Unlike stationary TBSs, HAPS platforms enable flexible deployment that can be dynamically optimized according to user distribution. Consequently, the HAPS deployment problem has been investigated using various analytical models across different application scenarios and architectural configurations.For example, Zong et al. [254] examined HAPS deployment for maximizing network transmission rate while meeting service requirements, where each GUE connects to the HAPS offering the best rate. A game-theoretic, self-organized optimization using a restricted spatial adaptive play algorithm was proposed and shown to converge in simulations. Although simplified compared to full SAGIN architectures, the results highlighted the potential of HAPS for self-organized deployment optimization.
@@
To meet 6G data rate demands, Cumali et al. [53] studied user selection in a downlink NOMA-based multi-user HAPS system. A joint user selection and correlation-based pairing algorithm was proposed to maximize beamforming gain, along with a polar-cap codebook design for limited-feedback HAPS channels under Rician fading. Simulations showed superior data rates over existing NOMA schemes. However, the model did not consider integration with terrestrial/satellite networks or support for aerial users, which are essential for 6G.
@@
While satellite deployment is less flexible than HAPS, its inherently wide coverage area provides opportunities for optimizing user association strategies, especially within multi-tier integrated architecture.Accordingly, the user association problem in the context of computational task offloading has been investigated by Zhang et al. [245]. The authors considered a disaster scenario where TBSs have been destroyed, and two-tier UEs, including both GUEs and UAVs, with limited computational capabilities offload disaster information to HAPS and/or LEO satellite servers for further processing. The study aimed to jointly optimize uplink fronthaul capacity and computation performance by modeling the problem as a weighted 3-uniform hypergraph, where each hyperedge weight represents the combined fronthaul uplink and computing rates. A local search-based algorithm was developed to find a near-optimal 3D hypergraph matching with maximum total weight. The model considered a LEO satellite with an enhanced MEC server providing backhaul to multiple HAPS, which cover disaster areas and offload computation, while UEs use orthogonal sub-channels within shared Ka-band RBs, avoiding uplink interference. In simulations with 6 HAPS and 20 UAVs over a 0.25 km² disaster area, the proposed hypergraph search algorithm achieved significant system sum rate gains over greedy search, showing strong practical potential.
@@
Later, a vHetNet comprising one GEO satellite, one HAPS, and several TBSs was examined by Liu, Dahrouj and Alouini [139]. In this setup, both the HAPS and TBSs are equipped with multiple antennas, enabling them to simultaneously serve multiple GUEs through RF links. The connection between the satellite and the HAPS is established via an FSO link. In this scenario, when a GUE is served by the HAPS, the required data is first transmitted from the satellite to the HAPS, and then relayed to the GUE.The system model employs a space division multiplexing scheme, where all RF links share the same central frequency. Consequently, the performance of the proposed architecture is highly dependent on the intra-cell and inter-cell interference among TBSs, intra-cell interference within the HAPS, and the interference between HAPS and TBSs. The study formulated a joint optimization problem for associating GUEs with either TBSs or the HAPS and designing corresponding beamforming vectors, aiming to maximize sum-rate under constraints of HAPS payload connectivity, FSO backhaul capacity, and power limits of HAPS/TBSs. An iterative algorithm optimized parameters step by step, and simulations demonstrated that integrated satellite-HAPS-terrestrial networks can significantly enhance the sum-rate.
@@
As heterogeneous integrated networks face load imbalance from uneven user distribution and mobility, Tu et al. [215] proposed a priority-based load-balancing scheme for a two-layer downlink SAGIN using network slicing to mitigate underutilization and overload across cells. The proposed two-layer radio access network (RAN) includes TBSs in the terrestrial layer (priority for user association) and UAVs/LEO satellites in the non-terrestrial layer. Overloaded TBSs offload traffic to the non-terrestrial layer via RAN slicing into high-throughput, low-delay, and wide-coverage slices, with lower-priority slices preferentially offloaded. The optimization problem jointly considered throughput, delay, coverage, user association, resource allocation, and UAV deployment. Results showed improved slice throughput and average delay. While HAPS were not included, the study tackled the critical load-balancing challenge in SAGIN. Prior to that, Shahid et al. [192] explored load balancing for 5G integrated satellite-terrestrial networks using multiple radio access technologys. The approach handled both intra- and inter-RAT UE offloading from overloaded cells by introducing the radio resource utilization ratio as a common load metric, enabling efficient traffic redistribution without overloading target cells. Simulations showed improved throughput and QoS. While limited to a GEO satellite, the metric can be generalized to other NTN components, including HAPS. Alam et al. [19] studied load balancing as a joint radio resource management problem—covering bandwidth allocation, UE association, and power control—in an integrated TN-NTN rural downlink network with multiple macro TBSs and LEO satellites. The proposed framework maximized the sum of log throughput across all UEs, improving overall network capacity. The framework tackled the optimization sequentially—first user association and bandwidth allocation under fixed power, then transmit power with the first two fixed. Simulations showed notable throughput gains over the maximum reference signal received power (RSRP) baseline and confirmed that integrating TNs and NTNs in rural areas can greatly reduce out-of-coverage UEs. In their subsequent study [18], the authors proposed a framework for bandwidth allocation, user association, and power control that can balance network capacity and energy consumption based on the current traffic conditions. By adapting the network behavior according to traffic demands, this approach significantly enhances mean network throughput during high-traffic periods and reduces power consumption during low-traffic periods.
@@
Based on the reviewed studies, we can conclude that user association in SAGINs that incorporate HAPS-mounted BSs and are fully integrated with TNs, still requires further research. It is particularly important to consider scenarios involving a realistically sufficient number of mobile ground and aerial UEs. Additionally, there is a need to investigate the user association challenge under the practical constraints imposed by interference management solutions and load balancing.
VI-D Multi-Link Connectivity
As mentioned earlier, some emerging 6G applications, such as aerial vehicle control and autonomous driving, impose stringent requirements in terms of both latency and connection reliability, a paradigm referred to as URLLC. Meanwhile, the Stage-2 architectural framework for MC across E-UTRA and NR is defined by 3GPP TS 37.340 [7]. It outlines the functions of the master node and secondary node, bearer splitting and duplication mechanisms, and the related signaling architecture that permits simultaneous connectivity to multiple radio nodes. In order to make clear how these processes might be used to enhance service continuity and resilience in HAPS-enabled NTN deployments, we have broadened the conversation.